### v1

@@ 7 Convergence Analysis in High Probability
    If $12r+s<$ 1, choose $\theta_{3}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{4}\geq\frac{2r}{(3+2r+s)\min\{r,1\}}=\frac{2r}{(3+2r+s)r}$, then
@@
We define the mapping $W_{0}:\mathcal{H}_{0}\rightarrow\mathcal{B}_{0}$ by $\ \sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\mapsto\sum_{i=1}^{n}\alpha_{i}W^{1/2}y_{i}\otimes\phi(x_{i})$ for any $n\in\mathbb{N}$, $x_{1},\cdots x_{n}\in\mathcal{X}$ and $y_{1},\cdots y_{n}\in\mathcal{Y}$. We see that $W_{0}$ is well-defined and linear. Moreover, for any $\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}=\sum_{i=1}^{n}\alpha_{i}\mathcal{K}(x_{i},\cdot)Wy_{i}\in\mathcal{H}_{0}$, we have
@@
$\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}(T+t_{0})^{-(1-\theta_{1})v}\begin{cases}\frac{1}{p+1}(T+t_{0})^{p+1},&\text{ when }p>-1,\\ \log(T+t_{0}),&\text{ when }p=-1,\\ \frac{t_{0}^{p+1}}{-1-p}\leq\frac{1}{-1-p},&\text{ when }p<-1,\end{cases}$
$\leq\delta^{\prime}\begin{cases}(T+t_{0})^{-(1-\theta_{1})v+p+1},&\text{ when }p>-1,\\ (T+t_{0})^{-(1-\theta_{1})v}\log(T+t_{0}),&\text{ when }p=-1,\\ (T+t_{0})^{-(1-\theta_{1})v},&\text{ when }p<-1,\end{cases}$
@@
[…]
@@
[…]
@@
[…]
@@
$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}\begin{cases}\frac{v}{v-1},&\text{ when }v>1,\\ \left(2-\theta_{1}\right)\log(T+t_{0}+1),&\text{ when }v=1,\\ \frac{\left(1-2^{\theta_{1}-1}\right)^{1-v}}{1-v}(T+t_{0}+1)^{(1-\theta_{1})(1-v)},&\text{ when }v<1,\end{cases}$
$=\widetilde{\delta^{\prime\prime}}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
@@
$\leq$ $2^{\bar{\eta}\bar{\lambda}}(\delta^{\prime}+\delta^{\prime\prime})\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1.\end{cases}$

### v3

@@ 7 Convergence Analysis in High Probability
    If $2r+s<1$, choose $\theta_{3}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{4}\geq\frac{2r}{(3+2r+s)\min\{r,1\}}=\frac{2r}{(3+2r+s)r}$, then
@@
[…]
@@
$\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}(T+t_{0})^{-(1-\theta_{1})v}\int_{0}^{\frac{T}{2}+1}(x+t_{0})^{p}dx$
@@
[…]
@@
$\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}\leq\int_{t+1}^{t+2}\frac{(u+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}\right]^{v}}du.$
@@
$\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}\leq 3^{-p}\int_{t+1}^{t+2}\frac{(u+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}\right]^{v}}du.$
@@
$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}\int_{0}^{(T+t_{0}+1)^{1-\theta_{1}}-(T/2+t_{0}+1)^{1-\theta_{1}}}\frac{(u+t_{0})^{p+\theta_{1}}}{1+\xi^{v}}d\xi$
$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}\int_{0}^{(T+t_{0}+1)^{1-\theta_{1}}(1-2^{\theta_{1}-1})}\frac{(u+t_{0})^{p+\theta_{1}}}{1+\xi^{v}}d\xi.$
@@
$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}\left(1+\int_{1}^{(T+t_{0}+1)^{1-\theta_{1}}(1-2^{\theta_{1}-1})}\frac{1}{\xi^{v}}d\xi\right)$
@@
[…]
@@
[…]